OpenStudy (anonymous):

f(x)=√(x+6)-x-5 f(√2+5)=???

OpenStudy (anonymous):

\[f(\sqrt{2}+5)=\sqrt{(\sqrt{2}+5)+6}-(\sqrt{2}+5)-5\]

OpenStudy (anonymous):

\[f(\sqrt{2}+5)=\sqrt{\sqrt{2}+11}-\sqrt{2}\]